There are more than 300 species of hummingbirds in the world and about 10 species regularly breed in the United States. One of which is the Ruby-throated Hmmingbird who can flap its wings with an angular speed of 324 rad/s. Calculate the period of its flapping motion.
Answer #1

As you know angular speed

W=2/T

Here W=324 rad/s

T=2 / W

T=2/324

T=0.0194 sec.